OpenStudy (anonymous):

Find the Derivative of G= ((5x^4)+(2x)-(1/4))/((x^3)-(cos(5)))

OpenStudy (experimentx):

20x^3 + 2 + 3x^2/4*1/((x^3)-(cos(5))^2

OpenStudy (anonymous):

Wow, that was really fast, I was going to get back on tomorrow. I appreciate it.

sam (.sam.):

@Headdesk Yes, we're really fast

OpenStudy (experimentx):

just a matter of practice ... derivative is not so difficult after all

OpenStudy (anonymous):

where is the 3x^2 coming from?

sam (.sam.):

Clearer form: \[\frac{-16 x^3 (5 \cos (5{}^{\circ})+1)-8 \cos (5{}^{\circ})+20 x^6+3 x^2}{4 \left(x^3-\cos (5{}^{\circ})\right)^2}\]

OpenStudy (anonymous):

I refuse to just take answers without learning something

sam (.sam.):

\[\text{Use the quotient rule, }\frac{d}{dx}\left(\frac{u}{v}\right)=\frac{v \frac{du}{dx}-u \frac{dv}{dx}}{v^2}\text{, where }u=5 x^4+2 x-\frac{1}{4}\text{ and }v=x^3-\cos (5{}^{\circ}):\]

sam (.sam.):

The page was not large enough \[v=x^3-\cos (5{}^{\circ})\]

OpenStudy (anonymous):

okay, I am going to write this down. Thanks again.